RC Meeting, Thursday, September 2, 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meeting at 10:00 am, (remote)

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Accelerator Status

  • Trying to deliver beam for physics to Hall B; issues last night with the beam size being too large and tails on vertical; Hall B accepted the beam for now but they want optimization ASAP;
  • ARC power supplies: they are finishing up ARC 8
  • DC power needs an access to look at PS of DOG8
  • Yan will check the injector setup in view of the Hall B beam issues
  • 5-pass setup: Sep. 5-6
  • Hall C setup: Sep. 7
  • Hall C start of Physics: September 8th

Hall Status

  • Target cooldown with He completed yesterday; H and D condensation will happen prior to first tune beam into the hall - probably on Monday
  • Steve L. rotated HMS and SHMS yesterday but not to the start-up angles of 7.5 deg (SHMS) and 12.5 deg (HMS); will double check with him that we can go ahead and rotate remotely; we should make sure SHMS and HMS are at the start-up angles before the hall sweep
  • Walter did the tech checklist and posted; all others: please do your checklists; we promised Ops that we would be ready for a sweep starting with late Friday afternoon

Notices in Effect for Shift Crew

Current Issues

  • Mismatch between the new FADC firmware and the FADC decoder/hcana being addresses by Mark
  • Interlock tripping on HMS Q1 [https://logbooks.jlab.org/entry/3896458} Steve took a look yesterday. Mark will exercise magnets today. --> need update
  • Target motion camera should be installed. Waiting to hear back from Joe B. --> need update
  • Checklist need to be done to lock up the Hall (detectors, beamline, magnets/roattion) --> sent out e-mail for checklists to be completed by Friday noon